OpenStudy (anonymous):

What are the solutions of x(x - 2) = 5?

OpenStudy (radar):

\[2\pm \sqrt{(-2)^{2}+20}\over 2\]\[2\pm \sqrt{24}\over2\]\[2\pm2\sqrt{6}\over2\]\[x=1\pm \sqrt{6}\]